abat9:

Which expression is equal to StartAbsoluteValue negative 12 EndAbsoluteValue minus StartAbsoluteValue Negative 3 EndAbsoluteValue? –15 –9 9 15

darkknight:

yo rewrite my g, the q is disoriented we can't read it

supie:

\(|−12|−|−3|\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight yo rewrite my g, the q is disoriented we can't read it \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

I-12I - I -3I = ?

Ari1234442:

|-12| - |-3| = 12 - 3 =

darkknight:

Alr ty supie, basically evaluate and solve throughtout, absolute valu;e of -12 is 12 absolute value of -3 is 3 so12-3 = ?

Ari1234442:

When there is an absolute value, anything within the absolute value becomes positive. Things outside the absolute values stay as they are.

Ari1234442:

what do you think the answer is